please help me solve the following word problem. tickets for a school dance cost $7 in advance or $11 at the door. If 109 students attend the dance and the total spent was $887, how many tickets were bought in advance and how many at the door?
Answer by mananth(16946)

x= Number of gate tickets
y= Number of tickets in advance

1 x + 1 y = 109 .............1
Total value
11 x + 7 y = 887 .............2
Eliminate y
multiply (1)by -7
Multiply (2) by 1
-7 x -7 y = -763
11 x + 7 y = 887
Add the two equations
4 x = 124
/ 4
x = 31
plug value of x in (1)
1 x + 1 y = 109
31 + y = 109
y = 109 -31
y = 78
y = 78
x= 31 Number of gate tickets
y= 78 Number of tickets in advance
m.ananth@hotmail.ca